# Study of hybrid stars with nonstrange quark matter cores

### v2

@@ III Structure of hybrid stars
The $M-\epsilon_{\rm c}$ relations are shown in Fig. 9. We can find that for stable neutron stars (whether they are hadron stars, quark stars, or hybrid stars), a larger $\epsilon_{\rm c}$ corresponds to a more massive star. For different values of $G_{1}$, the $\epsilon_{\rm pt}$ ($\epsilon_{\rm c}$) of the corresponding hybrid stars in the Panel (a) are around 670 (1000) MeV/fm3, and the difference is very small, which can also be seen in Table 2. In Panel (b), the difference of $\epsilon_{\rm pt}$ ($\epsilon_{\rm c}$) caused by different values of $B^{1/4}$ is more obvious than that in Panel (a).
@@
Figure 9: The $M-\epsilon_{\rm c}$ relations of quark stars and hybrid stars for (a) $B^{1/4}=153$ MeV and (b) $G_{1}=1.935$ GeV-2. The $M-\epsilon_{\rm c}$ relation of the hadron star described by BSK21 is shown by the red line. QS1935, QS1950, QS1972 (HS1935, HS1950, HS1972) refer to quark stars (hybrid stars) with $G_{1}=1.935,1.950,1.972$ GeV-2, respectively, and QS150, QS153, QS161 (HS150, HS153, HS161) refer to quark stars (hybrid stars) with $B^{1/4}=150,153,161$ MeV, respectively.
@@
We have also calculated the tidal deformability of hadron stars, quark stars and hybrid stars in this work, which is defined as Hinderer et al. (2010),
@@
where $\epsilon_{0}$ is the energy density at the surface of the star. The dimensionless parameter y can be calculated by solving the differential equations Hinderer et al. (2010),
@@
The $\Lambda-M$ relation is shown in Fig. 10. We can see that for quark stars and hadron stars described by the modified NJL model and BSK21 hadronic model, the corresponding values of $\Lambda(1.4M_{\odot})$ satisfy the constraint from GW170817, i.e., $\Lambda(1.4M_{\odot})=190_{-120}^{+390}$ Abbott and et al. (2018), except for the quark star with $G_{1}=1.935$ GeV-2 and $B^{1/4}=161$ MeV. For stable hybrid stars whose maximum masses are higher than 2 $M_{\odot}$ in this work, the corresponding hybrid EOSs demonstrate that neutron stars with the masses lower than 1.98 $M_{\odot}$ are still hadron stars and the quark matter cores do not exist inside them. Therefore, the $\Lambda-M$ relations from these hybrid EOSs are the same with that of hadron stars when $M\leq 1.98M_{\odot}$, and we do not show $\Lambda-M$ relations of hybrid stars in Fig. 10. In other words, according to the hybrid EOSs constrained in this work, the BNS in GW170817 whose masses are estimated to be 1.17-1.36 and 1.36-1.60 $M_{\odot}$ Abbott and et al. (2017a), respectively, should both be hadron stars.
@@
Figure 10: The $\Lambda-M$ relation of quark stars and hadron stars. QS1935, QS1950, QS1972 refer to quark stars with $B^{1/4}=153$ MeV, and $G_{1}=1.935,1.950,1.972$ GeV-2, respectively, and QS150, QS153, QS161 refer to quark stars with $G_{1}=1.935$ GeV-2, and $B^{1/4}=150,153,161$ MeV, respectively. The observational constraint from GW170817 ($\Lambda(1.4M_{\odot})=190_{-120}^{+390}$ Abbott and et al. (2018)) is also plotted for comparison.
@@
For the sake of completeness, the $M-R$ properties of hybrid stars constructed by nine representative hybrid EOSs are presented in Table 3, where ($R_{pt}$, $M_{pt}$) is related to the hadron-quark phase transition point, referring to the radius and mass of the most massive hadron star constructed by the hybrid EOS, and ($R_{\rm max}$, $M_{\rm max}$) is the radius and mass of the most massive hybrid star with a quark matter core. We can see that the masses of quark matter cores in these nine representative hybrid EOSs are in a range of 0.014-0.026 $M_{\odot}$. When $B^{1/4}=153$ MeV, a larger $G_{1}$ corresponds to a larger $M_{\rm max}$. For the same $G_{1}$, a larger $B^{1/4}$ leads to a larger $M_{\rm pt}$ and $M_{\rm max}$.
Table 3: The corresponding ($R_{pt}$, $M_{pt}$) and ($R_{\rm max}$, $M_{\rm max}$) points of nine representative hybrid EOSs.
$\quad\,G_{1}$ $\quad B^{1/4}$ $\,\,\,\,(R_{pt},M_{pt})$ $(R_{\rm max},M_{\rm max})$
$[{\rm GeV}^{-2}]$ $\,\,\,[{\rm MeV}]$ $\,\,\,\,\,[km,M_{\odot}]$ $\,\,\,\,\,\,[km,M_{\odot}]$
150150$\quad$ 150150 $(12.33,1.979)$ $\,(12.21,2.002)$
    1.935 153153$\quad$ 153153 $(12.28,2.020)$ $\,(12.20,2.038)$
161161$\quad$ 161161 $(12.15,2.094)$ $\,(12.11,2.108)$
149149$\quad$ 149149 $(12.33,1.979)$ $\,(12.22,2.002)$
    1.950 153153$\quad$ 153153 $(12.28,2.020)$ $\,(12.23,2.045)$
158158$\quad$ 158158 $(12.18,2.076)$ $\,(12.14,2.093)$
148148$\quad$ 148148 $(12.33,1.979)$ $\,(12.21,2.005)$
    1.972 153153$\quad$ 153153 $(12.25,2.039)$ $\,(12.21,2.059)$
157157$\quad$ 157157 $(12.18,2.076)$ $\,(12.14,2.092)$
IV Summary and discussion
@@
In this study, the modified 2-flavor NJL model and the BSK21 parametrization of the Skyrme interaction are introduced to investigate the nonstrange quark matter and hadronic matter in hybrid stars in light of a hypothesis that the quark matter may not be strange. To construct hybrid EOSs, the first-order hadron-quark phase transition and the corresponding Maxwell construction are considered.
@@
When the current quark mass is chosen as $m=3.5$ MeV, which is estimated by the particle data group, the modification of the coupling constant $G$ in the normal NJL model is helpful, because it is not only consistent with the QCD requirement in essence, but also in agreement with the lattice simulation results of $T_{\rm c}$. In the 2-flavor case when $T_{\rm c}=173\pm 8$ MeV, our calculations indicate that the parameter space of $G_{1}$ is limited to a range of 1.935-1.972 GeV-2, and the corresponding value of $G_{2}$ is from -1.582 to -0.743 GeV-5, implying that the normal 2-flavor NJL model with the four-quark scalar interaction (corresponding to the case of $G_{2}=0$ in our modified NJL model) is inconsistent with the lattice simulation results. For hybrid EOSs, the influence of $G_{1}$ is very small, but the influence of $B^{1/4}$ is obvious.
@@
Considering astronomical observations and the stability of hybrid stars, the parameter $B^{1/4}$ is constrained to be 150-161, 149-158, 148-157 MeV when $G_{1}=1.935,1.950,1.972$ GeV-2, respectively. The quark EOSs constructed with the modified NJL model in this work is soft, and thus cannot satisfy the 2 $M_{\odot}$ constraint of neutron stars and the $M-R$ constraints from NICER missions. It is noted that in some previous studies, the quark matter cores may not exist in compact stars under the Maxwell construction Özel (2006); Hoyos et al. (2016); Qin et al. (2023), or the maximum mass of quark matter cores may be larger than 0.6 $M_{\odot}$ and the BNS in GW170817 can be hybrid stars Ayriyan et al. (2021); Blaschke et al. (2020); Ayriyan et al. (2019); Pfaff et al. (2022). However, with the modified 2-flavor NJL model in this work, the hybrid EOSs with first-order hadron-quark transitions are still in agreement with current neutron star astronomical observations, and pure nonstrange quark matter cores can exist in hybrid stars, possessing a relatively small mass of 0.014-0.026 $M_{\odot}$. According to the hybrid EOSs constrained in this work, the BNS in GW170817 whose masses are estimated to be 1.17-1.36 and 1.36-1.60 $M_{\odot}$ Abbott and et al. (2017a), respectively, may be hadron stars.
Acknowledgements.
@@
This work is supported in part by the national Key Program for Science and Technology Research Development (2023YFB3002500), the National Natural Science Foundation of China (under Grants No. 12005192, No. 12075213, and No. 12233002), the Project funded by China Postdoctoral Science Foundation (No. 2020M672255, No. 2020TQ0287), National SKA Program of China No. 2020SKA0120300, the National Key R&\&&D Program of China (2021YFA0718500), the Natural Science Foundation of Henan Province of China (No. 242300421375), the Natural Science Foundation for Distinguished Young Scholars of Henan Province under grant number 242300421046, the start-up funding from Zhengzhou University. Y.F.H also acknowledges the support from the Xinjiang Tianchi Program.

### v3

@@ 3 Structure of hybrid stars
The $M-\epsilon_{\rm c}$ relations are shown in Fig. 9. We can find that for stable neutron stars (whether they are hadron stars, quark stars, or hybrid stars), a larger $\epsilon_{\rm c}$ corresponds to a more massive star. The $\epsilon_{\rm pt}$ ($\epsilon_{\rm c}$) of the hybrid stars whose corresponding hybrid EOSs satisfy the constraint of the lattice simulation results on $T_{\rm pc}$ is in a range of $856-886$ ($1180-1358$) MeV/fm3, which can also be seen in Table 2. However, for the normal NJL model with $m=7.5$ MeV, the corresponding $\epsilon_{\rm pt}$ ($\epsilon_{\rm c}$) of hybrid stars is about 980 (1700) MeV/fm3, quite different from the results of the modified NJL model.
@@
Figure 9: The $M-\epsilon_{\rm c}$ relations of quark stars and hybrid stars for $T_{\rm pc}=181$ MeV (left panel) and $m=7.5$ MeV (right panel), respectively. Line styles are the same as in Fig. 7.
@@
We have also calculated the tidal deformability of hadron stars, quark stars and hybrid stars in this work, which is defined as (Hinderer et al., 2010),
@@
where $\epsilon_{0}$ is the energy density at the surface of the star. The dimensionless parameter $y$ can be calculated by solving the differential equations  (Hinderer et al., 2010),
@@
The $\Lambda-M$ relation is shown in Fig. 10. We can see that for quark stars and hadron stars described by the modified NJL model and APR hadronic model, the corresponding values of $\Lambda(1.4M_{\odot})$ satisfy the constraint from GW170817, i.e., $\Lambda(1.4M_{\odot})=190_{-120}^{+390}$  (Abbott & et al., 2018). For stable hybrid stars whose maximum masses are higher than 2 $M_{\odot}$ in this work, the corresponding hybrid EOSs demonstrate that neutron stars with the masses lower than 1.974 $M_{\odot}$ are still hadron stars and the quark matter cores do not exist inside them. Therefore, the $\Lambda-M$ relations from these hybrid EOSs are the same with that of hadron stars when $M\leq 1.974$ $M_{\odot}$, and we do not show $\Lambda-M$ relations of hybrid stars in Fig. 10. In other words, according to the hybrid EOSs constrained in this work, the BNS in GW170817 whose masses are estimated to be $1.17-1.36$ and $1.36-1.60$ $M_{\odot}$ (Abbott & et al., 2017a), respectively, should both be hadron stars.
@@
Figure 10: The $\Lambda-M$ relation of quark stars and hadron stars. Here, m58, m75, m95 refer to quark stars with $T_{\rm pc}=181$ MeV and $m=5.8$, 7.5, 9.5 MeV, respectively, and m75G5290, m75G4552, m75G4430 refer to quark stars with $m=7.5$ MeV and $G_{1}=5.290$, 4.552, 4.430 GeV-2, respectively. The observational constraint from GW170817, $\Lambda(1.4M_{\odot})=190_{-120}^{+390}$  (Abbott & et al., 2018) is also plotted for comparison.
@@
For the sake of completeness, the $M-R$ properties of hybrid stars constructed by five representative hybrid EOSs are presented in Table 3, where ($R_{pt}$, $M_{pt}$) is related to the hadron-quark phase transition point, referring to the radius and mass of the most massive hadron star constructed by the hybrid EOS, and ($R_{\rm max}$, $M_{\rm max}$) is the radius and mass of the most massive hybrid star with a quark matter core. We can see that the masses of quark matter cores are in a range of $0.026-0.04$ $M_{\odot}$ in the framework of the modified NJL model. For a fixed $T_{\rm pc}$ ($m$), a larger $m$ ($G_{1}$) leads to a larger $M_{\rm max}$. Due to the constraint of $T_{\rm pc}$ from the lattice simulation results, $M_{\rm max}$ is about 0.1 $M_{\odot}$ lower than that of normal NJL model, while the quark matter core is about 0.015 $M_{\odot}$ heavier than that of normal NJL model.
Table 3: The corresponding ($R_{pt}$, $M_{pt}$) and ($R_{\rm max}$, $M_{\rm max}$) points of five representative hybrid EOSs.
$\quad\,m$ $\quad\,\,G_{1}$ $\,\,\,\,(R_{pt},M_{pt})$ $(R_{\rm max},M_{\rm max})$
$\,\,[{\rm MeV}]$ $[{\rm GeV}^{-2}]$ $\,\,\,\,\,[km,M_{\odot}]$ $\,\,\,\,\,\,[km,M_{\odot}]$
   5.8     3.4853.485\,\,\,\,3.4853.485 $(11.11,1.974)$ $\,(10.99,2.003)$
    4.4304.430\,\,\,\,4.4304.430 $(11.11,1.974)$ $\,(11.07,2.005)$
   7.5     4.5524.552\,\,\,\,4.5524.552 $(11.11,1.974)$ $\,(11.02,2.014)$
    5.2905.290\,\,\,\,5.2905.290 $(10.89,2.084)$ $\,(10.85,2.103)$
   9.5     5.5725.572\,\,\,\,5.5725.572 $(11.05,2.006)$ $\,(10.99,2.032)$
4 Summary
@@
In this study, the modified 2-flavor NJL model and the APR EOS with $A18+\delta\nu+UIX^{\ast}$ interaction are introduced to investigate the nonstrange quark matter and hadronic matter in hybrid stars in light of a hypothesis that the quark matter may not be strange. To construct hybrid EOSs, the first-order hadron-quark phase transition and the corresponding Maxwell construction are considered.
@@
It is noted that the modification of the coupling constant $G$ in the normal NJL model is helpful, because it is not only consistent with the QCD requirement in essence, but also in agreement with the lattice simulation results of $T_{\rm c}$. In the 2-flavor case, when $T_{\rm c}=173\pm 8$ MeV, the parameter space of $G_{1}$ can be constrained for a given current quark mass $m$, and the corresponding range of $G_{2}$ can also be determined. The results are shown in Table 1, implying that the normal 2-flavor NJL model with the four-quark scalar interaction (corresponding to the case of $G_{2}=0$ in our modified NJL model) is inconsistent with the lattice simulation results. For hybrid EOSs, the influence of $m$ and $G_{1}$ is very small when the constraint of $T_{\rm c}=173\pm 8$ MeV is taken into account. The hybrid EOSs derived from the modified NJL model are quite different from that of normal NJL model.
@@
Considering astronomical observations and the stability of hybrid stars, the parameter $m$ is constrained to be $5.8-9.5$ MeV, and the parameter space of $G_{1}$ can also be determined. The quark EOSs constructed with the modified NJL model in this work is soft, and thus cannot satisfy the 2 $M_{\odot}$ constraint of neutron stars and some $M-R$ constraints from NICER missions. It is noted that in some previous studies, the quark matter cores may not exist in compact stars under the Maxwell construction  (Özel, 2006; Hoyos et al., 2016; Qin et al., 2023), or the maximum mass of quark matter cores may be larger than 0.6 $M_{\odot}$ and the BNS in GW170817 can be hybrid stars  (Ayriyan et al., 2021; Blaschke et al., 2020; Ayriyan et al., 2019; Pfaff et al., 2022). However, with the modified 2-flavor NJL model in this work, the hybrid EOSs with first-order hadron-quark transitions are still in agreement with current neutron star astronomical observations, and pure nonstrange quark matter cores can exist in hybrid stars, possessing a relatively small mass of $0.026-0.04$ $M_{\odot}$. Due to the constraint of $T_{\rm c}=173\pm 8$ MeV, $M_{\rm max}$ of hybrid stars will be about 0.1 $M_{\odot}$ lower than that of normal NJL model. According to the hybrid EOSs constrained in this work, the BNS in GW170817 whose masses are estimated to be $1.17-1.36$ and $1.36-1.60$ $M_{\odot}$ (Abbott & et al., 2017a), respectively, may be hadron stars.
@@
We thank the anonymous referee for useful suggestions that led to an overall improvement of the presentation. This study is supported in part by the National Key Program for Science and Technology Research Development (2023YFB3002500), the National Natural Science Foundation of China (under Grants No. 12005192, No. 12075213, and No. 12233002), the Project funded by China Postdoctoral Science Foundation (No. 2020M672255, No. 2020TQ0287), National SKA Program of China No. 2020SKA0120300, the National Key R&\&&D Program of China (2021YFA0718500), the Natural Science Foundation of Henan Province of China (No. 242300421375), the Natural Science Foundation for Distinguished Young Scholars of Henan Province under grant number 242300421046, the start-up funding from Zhengzhou University. Y.F.H also acknowledges the support from the Xinjiang Tianchi Program.